OpenStudy (anonymous):

lim (x-2)/(x+2) x -3

OpenStudy (anonymous):

lim (x-2)/(x+2)= -3-2/-3+2=-5/-1=5 x -> -3

OpenStudy (anonymous):

thnx dude